OpenStudy (anonymous):

(8x + 70) - 150 = 8(x - 10)

OpenStudy (anonymous):

8x=70-150=8x-80 <==> 80=80

OpenStudy (anonymous):

So then, if we substitute for b, in y = mx + b.. b would equal 80?

OpenStudy (anonymous):

8x=70-150=8x-80 <==> -80=-80 ==> no solution here, what is the instruction here?

OpenStudy (anonymous):

can you post the whole question including any instruction?

OpenStudy (anonymous):

I'm supposed to be finding the slope of a line. And write an equation for it. With "Finding the line of best fit" Algebra 2

OpenStudy (anonymous):

I posted it.

OpenStudy (anonymous):

every line got x and y but you posted x's , can you check again?

OpenStudy (anonymous):

Exactly because i substituted y for the answer we got in the first post.

OpenStudy (anonymous):

confusing because I thought this a separte question